OpenStudy (anonymous):

The Law of Cosines lets you find missing measures in a triangle when you know the measures of two sides and the included angle, or three sides. True or false?

OpenStudy (ash2326):

True

OpenStudy (ash2326):

\[c^2=a^2+b^2-2ab \cos C\]